# Assessing the replicability of RCTs in RWE emulations

### v1

@@ 8.3 Availability of data and materials
The R package ReplicationSuccess, available on CRAN at https://CRAN.R-project.org/package=ReplicationSuccess, has been used for the computation of the sceptical $p$-value and all power calculations. The data can be found at https://gitlab.com/heyardr/hte-in-rwe/-/tree/main/data, and the code to reproduce all the Figures and Tables at https://github.com/CharlotteMicheloud/RWE.
@@ 8.6 Authors’ contributions
Jeanette Köppe: Conceptualization, Methodology, Formal analysis, Software, Writing - Original draft, Writing - Review & Editing, Visualization Charlotte Micheloud: Methodology, Formal analysis, Software, Writing - Original draft, Writing - Review & Editing, Visualization Stella Erdmann: Conceptualization, Methodology, Formal analysis, Review & Editing Rachel Heyard: Writing - Review & Editing Leonhard Held: Conceptualization, Methodology, Formal Analysis, Writing - Original draft, Writing - Review & Editing, Visualization

### v2

@@ 8.3 Availability of data and materials
The R package ReplicationSuccess, available on CRAN at https://CRAN.R-project.org/package=ReplicationSuccess, has been used for the computation of the sceptical $p$-value and all power calculations. The data and the code to reproduce all the Figures and Tables can be found at https://github.com/CharlotteMicheloud/RWE.
@@ 8.6 Authors’ contributions
Jeanette Köppe: Conceptualization, Methodology, Formal analysis, Software, Writing - Original draft, Writing - Review & Editing, Visualization Charlotte Micheloud: Methodology, Formal analysis, Software, Writing - Original draft, Writing - Review & Editing, Visualization Stella Erdmann: Conceptualization, Methodology, Formal analysis, Writing - Review & Editing Rachel Heyard: Data curation, Writing - Review & Editing Leonhard Held: Conceptualization, Methodology, Formal Analysis, Writing - Original draft, Writing - Review & Editing, Visualization
@@ Appendix A Replication power
Appendix B Confidence intervals for combined effects
Study NI RCT pooled RWE meta analysis sceptical $p$-value
margin HR (95% CI) HR (95% CI) HR (95% CI) (one-sided 97.5% CI)
1 LEADER 1.30 0.87 (0.78; 0.97) 0.82 (0.76; 0.87) 0.83 (0.79; 0.88) (0; 0.91)
2 DECLARE 1.30 0.83 (0.73; 0.95) 0.69 (0.59; 0.81) 0.77 (0.70; 0.85) (0; 0.88)
3 EMPA-REG 1.30 0.86 (0.74; 0.99) 0.83 (0.73; 0.95) 0.84 (0.76; 0.93) (0; 0.92)
4 CANVAS 1.30 0.86 (0.75; 0.97) 0.77 (0.70; 0.85) 0.80 (0.74; 0.87) (0; 0.91)
5 CARMELINA 1.30 1.02 (0.89; 1.17) 0.90 (0.84; 0.96) 0.92 (0.87; 0.98) (0; 1.07)
6 TECOS 1.30 0.98 (0.88; 1.09) 0.89 (0.86; 0.91) 0.90 (0.87; 0.92) (0; 1.01)
7 SAVOR-TIMI 1.30 1.00 (0.89; 1.12) 0.81 (0.76; 0.86) 0.85 (0.80; 0.90) (0; 1.04)
8 TRITON-TIMI 1.00 0.81 (0.73; 0.90) 0.88 (0.79; 0.97) 0.84 (0.79; 0.91) (0; 0.92)
9 PLATO 1.00 0.84 (0.77; 0.92) 0.92 (0.83; 1.02) 0.87 (0.82; 0.93) (0; 0.96)
10 ARISTOTLE 1.38 0.79 (0.66; 0.95) 0.68 (0.61; 0.76) 0.71 (0.64; 0.78) (0; 0.85)
11 RE-LY 1.46 0.66 (0.53; 0.82) 0.73 (0.60; 0.90) 0.70 (0.60; 0.81) (0; 0.80)
12 ROCKET-AF 1.46 0.79 (0.66; 0.96) 0.70 (0.62; 0.80) 0.73 (0.65; 0.81) (0; 0.85)
13 EINSTEIN-DVT 2.00 0.68 (0.44; 1.04) 0.75 (0.62; 0.90) 0.74 (0.62; 0.88) (0; 0.86)
14 EINSTEIN-PE 2.00 1.12 (0.75; 1.64) 0.67 (0.55; 0.80) 0.74 (0.62; 0.87) (0; 1.29)
15 RECOVER II 2.75 1.08 (0.64; 1.80) 1.15 (0.74; 1.78) 1.12 (0.80; 1.57) (0; 1.48)
16 AMPLIFY 1.80 0.84 (0.60; 1.18) 0.81 (0.54; 1.23) 0.83 (0.64; 1.07) (0; 1.03)
17 RECORD1 1.95 0.25 (0.14; 0.47) 0.17 (0.10; 0.29) 0.20 (0.13; 0.30) (0; 0.33)
18 TRANSCEND 1.00 0.92 (0.81; 1.05) 0.88 (0.81; 0.96) 0.89 (0.83; 0.96) (0; 0.97)
19 ON-TARGET 1.13 1.01 (0.94; 1.09) 0.83 (0.77; 0.90) 0.92 (0.87; 0.97) (0; 1.04)
20 HORIZON-PIVOTAL 1.00 0.59 (0.42; 0.83) 0.75 (0.58; 0.97) 0.69 (0.56; 0.84) (0; 0.84)
21 DAPA-CKD 1.00 0.61 (0.51; 0.72) 0.80 (0.52; 1.26) 0.63 (0.54; 0.74) (0; 0.99)
22 PARADIGM-HF 1.00 0.80 (0.73; 0.87) 1.02 (0.91; 1.14) 0.88 (0.82; 0.94) (0; 1.07)
23 P04334 1.00 0.56 (0.44; 0.72) 0.78 (0.62; 0.97) 0.67 (0.57; 0.79) (0; 0.86)
24 D5896 2.00 1.07 (0.70; 1.65) 1.38 (0.90; 2.13) 1.21 (0.90; 1.65) (0; 1.68)
25 IMPACT 1.00 0.85 (0.80; 0.90) 1.13 (1.04; 1.23) 0.93 (0.89; 0.98) (0; 1.17)
26 POET-COPD 1.00 0.83 (0.77; 0.90) 1.02 (0.93; 1.12) 0.90 (0.85; 0.96) (0; 1.06)
27 INSPIRE 1.00 0.97 (0.84; 1.12) 0.93 (0.90; 0.96) 0.93 (0.90; 0.96) (0; 1.01)
28 CAROLINA 1.30 0.98 (0.84; 1.14) 0.91 (0.79; 1.05) 0.94 (0.85; 1.05) (0; 1.05)
29 PRONOUNCE 1.00 1.28 (0.59; 2.79) 1.35 (0.94, 1.93) 1.34 (0.96; 1.85) (0; 1.85)
Table B.1: Confidence intervals (CIs) for the hazard ratio (HR) from each RCT and pooled RWE. Combined confidence intervals based on fixed effect meta-analysis and sceptical $p$-value are also reported, as well as the non-inferiority (NI) margin (1.00 for superiority studies).